# Shaving Weights with Occam’s Razor: Bayesian Sparsification for Neural Networks using the Marginal Likelihood

### v1

@@ Appendix A Proof for Diagonal Prior in a Kronecker-factored Eigenbasis
We prove this result in two steps. First, we show what the optimum looks like in terms of the Frobenius norm. Second, we show how to simplify the results to enable efficient computation without computing Kronecker products. We have a KFAC Hessian approximation ${\mathbf{A}}\otimes{\mathbf{B}}$ with ${\mathbf{A}}\in\mathbb{R}^{D_{\textrm{in}}\times D_{\textrm{in}}}$ and ${\mathbf{B}}\in\mathbb{R}^{D_{\textrm{out}}\times D_{\textrm{out}}}$ where the dimensionalities $D_{\cdot}$ depend on the layer type (Martens and Grosse, 2015). In the case of a fully-connected layer, these are simply the dimensionality of the in- and output hidden representation. The same layer will have $D_{\textrm{in}}\times D_{\textrm{out}}$ parameters and thus the corresponding diagonal prior precision is given by ${\boldsymbol{\delta}}\in\mathbb{R}^{D_{\textrm{in}}D_{\textrm{out}}}$. For the Laplace approximation, the eigendecomposition of individual Kronecker factors is already computed as ${\mathbf{A}}={\mathbf{Q}}_{A}{\boldsymbol{\Lambda}}_{A}{\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}$ and similarly for ${\mathbf{G}}$ as shown in Equation 5. Recall also that $\operatorname{diag}(\cdot)$ turns a vector into a diagonal matrix and extracts the diagonal entries of a matrix into a vector. We are interested in the Frobenius-optimal diagonal perturbation of the eigenvalues so as to maintain the efficiency structure of the KFAC and thus the downstream Laplace approximation:
$\operatorname*{arg\,min}_{\hat{{\boldsymbol{\delta}}}}\|({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G})({\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G}+\operatorname{diag}{(\hat{{\boldsymbol{\delta}}})})({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})-({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G})({\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G})({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})+\operatorname{diag}{({\boldsymbol{\delta}})}\|_{F}^{2}$
$=$ $\operatorname*{arg\,min}_{\hat{{\boldsymbol{\delta}}}}\|{\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G}+\operatorname{diag}{(\hat{{\boldsymbol{\delta}}})}-{\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G}+({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})\operatorname{diag}{({\boldsymbol{\delta}})}({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G})\|_{F}^{2}$
$=$ $\operatorname{diag}{(({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})\operatorname{diag}{({\boldsymbol{\delta}})}({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G}))},$
@@
$\operatorname{diag}{({\mathbf{Q}}^{\mathrm{\textsf{T}}}\operatorname{diag}{({\mathbf{d}})}{\mathbf{Q}})}=({\mathbf{Q}}^{\mathrm{\textsf{T}}}\circ{\mathbf{Q}}^{\mathrm{\textsf{T}}}){\mathbf{d}},$
@@
$\operatorname{diag}(({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})\operatorname{diag}{({\boldsymbol{\delta}})}({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G}))$ $=(({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})\circ({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})){\boldsymbol{\delta}}$
$=(({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\circ{\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}})\otimes({\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}}\circ{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})){\boldsymbol{\delta}}$
$=\operatorname{vec}(({\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}}\circ{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})\operatorname{mat}({\boldsymbol{\delta}})({\mathbf{Q}}_{A}\circ{\mathbf{Q}}_{A}))=\operatorname{vec}({\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})^{2}\operatorname{mat}({\boldsymbol{\delta}}){\mathbf{Q}}_{A}^{2},$
@@ Appendix B Additional Results
B.1 Baselines
Figure 7 illustrates that both MAP and SPAM achieve similar levels of performance throughout the training process. This observation underscores that SPAM’s enhanced sparsifiability is achieved without compromising the unpruned performance. Furthermore, the comparable unpruned accuracies of SPAM and MAP models indicate that SPAM’s sparsifiability benefits are not merely a result of higher baseline accuracies, but rather a distinct advantage offered by the SPAM methodology. The sparsification methods are performed on these models in a way that once the model is trained for a specific seed, we copy it and use it to perform the different sparsification methods, we repeat the steps for a minimum of 4 different seeds ensuring the robustness of our findings.
(a) FC on Cancer
(b) FC on MNIST
(c) LeNet on MNIST
(d) LeNet on FashionMNIST
(e) ResNet on CIFAR10 (Not Augmented)
(f) MLPMixer
Figure 7: Training curves for MAP and SpaM training with different priors and Hessian approximations. We see that all methods achieve a similar performance by the end of training.
@@ B.2 Tables
In tables 2 and 3, we present our results comparing different methods using MAP and SpaM with various priors. Notably, SpaM with Diag LA and parameter-wise priors significantly outperforms MAP and other SpaM variants at high sparsity levels.
@@
Table 2: Comparison of pruning accuracies of SpaM training with different pruning criteria, Hessian approximations, and priors for post-hoc pruning LeNet on MNIST.
@@
Table 3: Comparison of pruning accuracies of SpaM training with different pruning criteria, Hessian approximations, and priors for post-hoc pruning MLP-Mixer (2 blocks) on MNIST.
Table 4: NLL of unstructured pruned ResNets on CIFAR-10. The best training method for each pruning criterion is highlighted in green, showing that SpaM improves performance over MAP for most criteria. The best performances (lowest NLL) overall at each sparsity level are shown in bold, showing that our OPD criterion outperforms the others at most sparsity levels.
@@ B.3 Prior effects
Figure 8 and Figure 9 illustrate our findings when applying SpaM with various priors for both OPD and GraSP. Notably, Diag LA, using parameter-wise priors, excels in high sparsity scenarios, even with complex models and datasets like ResNets. Furthermore, for MLPmixer, we observe that SpaM variants, employing parameter-wise priors and layerwise approaches, preserve baseline accuracy even at extreme sparsities of 99%.
(a) FC on Cancer
(b) FC on MNIST
(c) LeNet on MNIST
(d) LeNet on FashionMNIST
(e) ResNet on CIFAR10 (Not Augmented)
(f) MLPMixer
Figure 8: Effect of different priors and Hessian approximations on the sparsification performance with SpaM-OPD. The diagonal approximation with parameter-wise priors is a strong choice, especially at higher sparsities, while the KFAC approximation with layerwise prior yields slightly better performances at lower sparsities.
(a) FC on Cancer
(b) FC on MNIST
(c) LeNet on MNIST
(d) LeNet on FashionMNIST
(e) ResNet on CIFAR10 (Not Augmented)
(f) MLPMixer
Figure 9: Priors and Hessian approximations for GraSP pruning. We see that the effects are qualitatively similar to pruning with OPD.
@@ B.4 Unit-wise and Parameter-wise KFAC for GraSP
As shown in Section 5, networks trained using SpaM and parameter-wise priors were able to maintain a high accuracy at challenging sparsity levels up to 99%. Moreover, parameter-wise KFAC and unit-wise priors showed high performance for the OPD pruning approach. We show in Figure 10 that the combination of SpaM and these priors leverage the performance of methods like GraSP.
Figure 10: Priors and Hessian approximations for GraSP pruning with SpaM. We see that the effects are qualitatively similar to pruning with OPD.
@@ B.5 One Shot Efficiency
As seen in Figure 11, our proposed post-hoc pruning criterion, OPD, consistently demonstrates stable performance across diverse model architectures and datasets, achieving significant sparsity levels without the need for fine-tuning. It seamlessly operates either post-training or with pre-trained models, providing a highly flexible and versatile solution.
@@
Figure 11: SpaM post-hoc pruning efficiency with optional fine-tuning after the pruning. Unlike other pruning criteria, OPD does not require additional tuning to achieve optimal performance across different architectures and often still outperforms the other fine-tuned methods.
B.6 Modern Architectures
B.6.1 Wide ResNet
@@
In Figure 12, we demonstrate how SpaM enhances the sparsity performance of Wide ResNet models. This is specifically illustrated in the case of OPD, GraSP, and Magnitude, all while maintaining a low Brier score, ECE, and NLL up to 95% sparsity.
@@
Figure 12: SpaM post-hoc efficiency for Wide ResNet on CIFAR10. Leveraging OPD, GraSP, and Magnitude performance under SpaM in comparison to MAP show superior test accuracy at increased sparsity levels coupled with a low ECE, Brier Score, and NLL.
B.6.2 Vision Transformer
@@
Figure 13 demonstrates the impact of SpaM on Unstructured Pruning for a Vision Transformer (ViT) trained on MNIST. These results align with the findings presented in Section 5 with SpaM diagonal LA and parameter-wise priors leveraging the sparsifiability of models using OPD, Magnitude and GraSP, maintaining a test accuracy of 97% for OPD and Magnitude at 95% sparsity compared to an accuracy of lower than 20% under MAP for the same methods. This serves as a proof-of-concept for vision transformers but efficacy has to be verified at a larger scale where such models perform best.
@@
Figure 13: SpaM post-hoc efficiency for ViT on MNIST. Leveraging OPD, GraSP, and Magnitude performance under SpaM in comparison to MAP show superior test accuracy at increased sparsity levels coupled with a low ECE, Brier Score, and NLL.
B.7 Visualization of Pruning Process
To visualize the model’s structural evolution during pruning, we present a series of filter bank visualizations that capture the key stages of transformation, from the initial dense architecture to the final compact form (Figures 14, 15,16, 17).
ptRefer to captionConv1ptRefer to captionConv2ptRefer to captionConv3ptRefer to captionLin1ptRefer to captionLinoutLeNet Baseline
Figure 14: Visualization of model weights for unpruned LeNet on FashionMNIST.
ptRefer to captionConv1ptRefer to captionConv2ptRefer to captionConv3ptRefer to captionLin1ptRefer to captionLinoutLeNet on FashionMNIST 95% Sparse
Figure 15: Visualization of model weights for 95 % pruned LeNet on FashionMNIST. Black refers to pruned weights or a part of a filter. An entire black square refers to an entire filter being pruned. A black row or column represents a pruned neuron.
ptRefer to captionConv1ptRefer to captionConv2ptRefer to captionConv3ptRefer to captionLin1ptRefer to captionLinoutLeNet on FashionMNIST 95% Sparse
Figure 16: Visualization of model weights for 95 % pruned LeNet on FashionMNIST after the zeroing stage. Black refers to pruned weights or a part of a filter. An entire black square refers to an entire filter being pruned. A black row or column represents a pruned neuron. (Model size = 241 KB).
ptRefer to captionConv1ptRefer to captionConv2ptRefer to captionConv3ptRefer to captionLin1ptRefer to captionLinoutLeNet on FashionMNIST 95% Sparse
Figure 17: Visualization of model weights for 95 % pruned LeNet on FashionMNIST after the compression stage. Black refers to pruned weights or a part of a filter. An entire black square refers to an entire filter being pruned. A black row or column represents a pruned neuron. (Model size = 12 KB).
B.8 Network Compression

### v2

@@ Appendix A Proof for Diagonal Prior in a Kronecker-factored Eigenbasis
We prove this result in two steps. First, we show what the optimum looks like in terms of the Frobenius norm. Second, we show how to simplify the results to enable efficient computation without computing Kronecker products. We have a KFAC Hessian approximation ${\mathbf{A}}\otimes{\mathbf{B}}$ with ${\mathbf{A}}\in\mathbb{R}^{D_{\textrm{in}}\times D_{\textrm{in}}}$ and ${\mathbf{B}}\in\mathbb{R}^{D_{\textrm{out}}\times D_{\textrm{out}}}$ where the dimensionalities $D_{\cdot}$ depend on the layer type [18]. In the case of a fully-connected layer, these are simply the dimensionality of the in- and output hidden representation. The same layer will have $D_{\textrm{in}}\times D_{\textrm{out}}$ parameters and thus the corresponding diagonal prior precision is given by ${\boldsymbol{\delta}}\in\mathbb{R}^{D_{\textrm{in}}D_{\textrm{out}}}$. For the Laplace approximation, the eigendecomposition of individual Kronecker factors is already computed as ${\mathbf{A}}={\mathbf{Q}}_{A}{\boldsymbol{\Lambda}}_{A}{\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}$ and similarly for ${\mathbf{G}}$ as shown in Equation 5. Recall also that $\operatorname{diag}(\cdot)$ turns a vector into a diagonal matrix and extracts the diagonal entries of a matrix into a vector. We are interested in the Frobenius-optimal diagonal perturbation of the eigenvalues so as to maintain the efficiency structure of the KFAC, and, thus, the downstream Laplace approximation:
$\operatorname*{arg\,min}_{\hat{{\boldsymbol{\delta}}}}\|({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G})({\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G}+\operatorname{diag}{(\hat{{\boldsymbol{\delta}}})})({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})$
$\qquad\qquad-({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G})({\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G})({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})+\operatorname{diag}{({\boldsymbol{\delta}})}\|_{F}^{2}$
$=$ $\operatorname*{arg\,min}_{\hat{{\boldsymbol{\delta}}}}\|{\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G}+\operatorname{diag}{(\hat{{\boldsymbol{\delta}}})}-{\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G}+({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})\operatorname{diag}{({\boldsymbol{\delta}})}({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G})\|_{F}^{2}$
$=$ $\operatorname{diag}{(({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})\operatorname{diag}{({\boldsymbol{\delta}})}({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G}))},$
@@
$\operatorname{diag}{({\mathbf{Q}}^{\mathrm{\textsf{\tiny T}}}\operatorname{diag}{({\mathbf{d}})}{\mathbf{Q}})}=({\mathbf{Q}}^{\mathrm{\textsf{\tiny T}}}\circ{\mathbf{Q}}^{\mathrm{\textsf{\tiny T}}}){\mathbf{d}},$
@@
$\operatorname{diag}(({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})\operatorname{diag}{({\boldsymbol{\delta}})}({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G}))$ $=(({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})\circ({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})){\boldsymbol{\delta}}$
$=(({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\circ{\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}})\otimes({\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}}\circ{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})){\boldsymbol{\delta}}$
$=\operatorname{vec}(({\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}}\circ{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})\operatorname{mat}({\boldsymbol{\delta}})({\mathbf{Q}}_{A}\circ{\mathbf{Q}}_{A}))$
$=\operatorname{vec}({\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})^{2}\operatorname{mat}({\boldsymbol{\delta}}){\mathbf{Q}}_{A}^{2},$
@@ Appendix B Additional Results
B.1 Baseline training
Figure B1 illustrates that both MAP and SPAM achieve similar levels of performance throughout the training process. This observation underscores that SPAM’s enhanced sparsifiability is achieved without compromising the unpruned performance. Furthermore, the comparable unpruned accuracies of SPAM and MAP models indicate that SPAM’s sparsifiability benefits are not merely a result of higher baseline accuracies, but rather a distinct advantage offered by the SPAM methodology. The sparsification methods are performed on these models in a way that once the model is trained for a specific seed, we copy it and use it to perform the different sparsification methods; we repeat the steps for a minimum of 4 different seeds, ensuring the robustness of our findings. In addition, the model trained with SpaM only uses a single forward pass over the pruned architecture during inference, thus guaranteeing a fair comparison with the baselines. The posterior is solely used to estimate the marginal likelihood during SpaM training.
Figure B1: Training curves for MAP and SpaM training with different priors and Hessian approximations. We see that all methods achieve a similar performance by the end of training.
@@ B.2 Tables
In tables B1 and B2, we present our results comparing different methods using MAP and SpaM with various priors. Notably, SpaM with Diag LA and parameter-wise priors significantly outperforms MAP and other SpaM variants at high sparsity levels.
Table B1: Comparison of pruning accuracies of SpaM training with different pruning criteria, Hessian approximations, and priors for post-hoc pruning LeNet on MNIST.
@@
Table B2: Comparison of pruning accuracies of SpaM training with different pruning criteria, Hessian approximations, and priors for post-hoc pruning MLP-Mixer (2 blocks) on MNIST.
@@
Table B3: NLL of unstructured pruned ResNets on CIFAR-10. The best training method for each pruning criterion is highlighted in green, showing that SpaM improves performance over MAP for most criteria. The best performances (lowest NLL) overall at each sparsity level are shown in bold, showing that our OPD criterion outperforms the others at most sparsity levels.
@@ B.3 Prior effects
Figure B2 and Figure B3 illustrate our findings when applying SpaM with various priors for both OPD and GraSP. Notably, Diag LA, using parameter-wise priors, excels in high-sparsity scenarios, even with complex models and datasets like ResNets. Furthermore, for MLPmixer, we observe that SpaM variants, employing parameter-wise priors and layerwise approaches, preserve baseline accuracy even at extreme sparsities of 99%.
Figure B2: Effect of different priors and Hessian approximations on the sparsification performance with SpaM-OPD. The diagonal approximation with parameter-wise priors is a strong choice, especially at higher sparsities, while the KFAC approximation with layerwise prior yields slightly better performances at lower sparsities.
Figure B3: Priors and Hessian approximations for GraSP pruning with SpaM and MAP, including our newer priors (parameter-wise and unit-wise). We see that the effects are qualitatively similar to pruning with OPD.
@@ B.4 Unit-wise and Parameter-wise KFAC for GraSP
As shown in Section 5, networks trained using SpaM and parameter-wise priors were able to maintain a high accuracy at challenging sparsity levels up to 99%. Moreover, parameter-wise KFAC and unit-wise priors showed high performance for the OPD pruning approach. We show in Figure B3 that the combination of SpaM and these priors with the GraSP criterion yield qualitatively similar performance rankings as with our OPD criterion.
@@ B.5 One Shot Efficiency
As seen in Figure B4, our proposed post-hoc pruning criterion, OPD, consistently demonstrates stable performance across diverse model architectures and datasets, achieving significant sparsity levels without the need for fine-tuning. It seamlessly operates either post-training or with pre-trained models, providing a highly flexible and versatile solution.
@@
Figure B4: SpaM post-hoc pruning efficiency with optional fine-tuning after the pruning. Unlike other pruning criteria, OPD does not require additional tuning to achieve optimal performance across different architectures and often still outperforms the other fine-tuned methods.
B.6 Online pruning.
Figure B5: Predictive performance as a function of sparsity level in unstructured pruning. We include our online pruning approach that progressively prunes a model during the training compared to the other curves demonstrating the performance of 10 pruned models based on a converged baseline. our online pruning approach is often competitive with post-hoc pruning.
Figure B5 shows that our online version of SpaM, which uses the marginal likelihood and OPD during training to iteratively prune the network, reaches comparable performance levels to the post-hoc version, thus offering a computationally even more convenient way to effectively sparsify neural networks. In the online approach, we prune jointly during SpaM training to reach a target sparsity; we perform the sparsity updates (dynamic masking) based on the marginal likelihood optimization parameters we refer to as n_epochs_burnin and marglik_frequency. Here, n_epochs_burnin specifies when we start the marginal likelihood optimization and marglik_frequency specifies after how many epochs we update the estimate. Pruning occurs only after a new marginal likelihood calculation, and the sparsity percentage is adjusted incrementally to reach the target by the training’s end. The curve of OPD-Online reflects training progress (x-axis explicitly epochs / reached sparsity), which explains the curve of LeNet on CIFAR-10 that is still converging while pruning. At the same time, the one-shot post-training approach reflects converged models copied and pruned at different sparsity levels.
B.7 Comparing SpaM to L1 regularization
We conducted experiments using L1 regularization, which is well-known for inducing sparsity in neural networks. To optimize performance, we performed an extensive search for the appropriate L1 regularization coefficient and applied various strengths during training.
We found that SpaM consistently outperforms L1 regularization, achieving much higher levels of sparsity while maintaining the network’s predictive performance as shown in Figure B6. L1 regularization, while effective at inducing sparsity, often proved too aggressive, leading to networks that were excessively pruned, negatively affecting performance. In contrast, our method offers a key advantage: it adapts the regularization to each individual weight based on the data, rather than relying on a single global parameter, allowing for more nuanced control over sparsity.
Figure B6: Comparing SpaM and MAP to L1 Regularization. We see that SpaM consistently outperforms L1 regularization.
B.8 Additional Pruning Results
B.8.1 Wide ResNet
@@
In Figure B7, we demonstrate how SpaM enhances the sparsity performance of Wide ResNet models. This is specifically illustrated in the case of OPD, GraSP, and Magnitude, all while maintaining a low Brier score, ECE, and NLL up to 95% sparsity.
@@
Figure B7: SpaM post-hoc efficiency for Wide ResNet on CIFAR10. Leveraging OPD, GraSP, and Magnitude performance under SpaM in comparison to MAP show superior test accuracy at increased sparsity levels coupled with a low ECE, Brier Score, and NLL.
B.8.2 Vision Transformer
@@
Figure B8 demonstrates the impact of SpaM on Unstructured Pruning for a Vision Transformer (ViT) trained on MNIST. These results align with the findings presented in Section 5 with SpaM diagonal LA and parameter-wise priors leveraging the sparsifiability of models using OPD, Magnitude and GraSP, maintaining a test accuracy of 97% for OPD and Magnitude at 95% sparsity compared to an accuracy of lower than 20% under MAP for the same methods. This serves as a proof-of-concept for vision transformers but efficacy has to be verified at a larger scale where such models perform best.
@@
Figure B8: SpaM post-hoc efficiency for ViT on MNIST. Leveraging OPD, GraSP, and Magnitude performance under SpaM in comparison to MAP show superior test accuracy at increased sparsity levels coupled with a low ECE, Brier Score, and NLL.
B.8.3 GPT-2
We demonstrate the efficacy of OPD on a pre-trained GPT-2 model (124M parameters), fine-tuned for sentiment analysis on the IMDB dataset. To manage computational resources, we limit both the Laplace approximation and SpaM to two steps. Despite this constraint, OPD maintains high predictive performance even at 60% sparsity, as shown in Figure B9. This suggests that extending SpaM optimization with more epochs and a more refined posterior could further enhance performance.
Figure B9: GPT-2 (124M) on IMDB. We tune GPT-2 for sentiment analysis on IMDB datasets. Our results show that OPD maintains significantly higher accuracy than other methods, which degrade towards random classifier performance (50%) at 60% sparsities.
B.9 Visualization of the Pruning Process
In order to illustrate the structural evolution of the model throughout the different sparsification approaches (unstructured and structured), we provide a sequence of filter bank visualizations that delineate the principal stages of pruning, progressing from the initial dense architecture to the ultimate compact configuration. These visualizations also reveal the influence of parameter-wise and unit-wise priors on the weights (Figure B10).
Figure B10: Visualization of the sparsification effect on the model’s layers for LeNet on FashionMNIST sparsified at 95% using both the structured and unstructured approach. Blocks in black refer to masked filters, and columns refer to neurons pruned. We see that for unit-wise priors, a 95% sparsity yields more entire kernels and neurons being masked compared to parameter-wise priors.
B.10 Network Compression